OpenStudy (anonymous):

Write the percent as a decimal 4%, 23%, 1.5%, 2.5%

OpenStudy (anonymous):

Show work

OpenStudy (anonymous):

.04, .23, 150. 250.

OpenStudy (anonymous):

4/100 , 23/100 , 1.5/100,2.5/100 .04,.23,.015,.025

OpenStudy (anonymous):

just move the decimal two places to the left on each of them.

OpenStudy (anonymous):

4/100=0.04 23/100=0.23 1.5/100=0.015 2.5/100=0.025

OpenStudy (anonymous):

oopps messed up my decimals the last two should be .025 ad .015